OpenStudy (anonymous):

Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

@Mandy86 help

OpenStudy (anonymous):

O crap!

OpenStudy (anonymous):

ok so you can get rid of the x's to get y

OpenStudy (anonymous):

show work

OpenStudy (anonymous):

3y=11 y=9

OpenStudy (anonymous):

3y − 2x = 11 y + 2x = 9 ---------------add both equations 4y = 20 --------divide both sides by 4 y = 5 plug that into any of the equations y + 2x = 9 5 + 2x = 9 subtract 5 2x = 9- 5 =4 divide by 2 x = 2

OpenStudy (anonymous):

that doesnt look right hold on im gonna write the whole question

OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9 I have done part 1

OpenStudy (anonymous):

i think its like a graphing equation so i have to get an answer, for example like (3,2)

OpenStudy (anonymous):

ok write it down in a point form

OpenStudy (anonymous):

not sure how to

OpenStudy (anonymous):

i know y=mx+b

OpenStudy (anonymous):

(x,y) what are the values we found?

OpenStudy (anonymous):

i mean i found

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

oh wait nvm i see it now i just didnt see the Y

OpenStudy (anonymous):

so (5,2)

OpenStudy (anonymous):

wrong (x,y) not (y,x)

OpenStudy (anonymous):

2,5

OpenStudy (anonymous):

correct